OpenStudy (s):

Average velocity question :A car makes a trip due north for three-fourths of the time and due south one-fourth of the time. The average northward velocity has a magnitude of 46 m/s, and the average southward velocity has a magnitude of 32 m/s. Taking northward to be the positive direction, what is the average velocity for the trip?

OpenStudy (shane_b):

You can do this one a few different ways but I think the simplest is to just calculate the average of the values: \[\frac{46+46+46-32}{4}=?\]

OpenStudy (perl):

average velocity = total distance / total time Let the total time be t average velocity = ( 46 * 3/4 * t + (-32)*(1/4*t) ) / ( 3/4*t + 1/4*t) = 46*3/4 - 32*1/4